Answer the following problems


1. Consider a given bond that has five years maturity, Br.1000 face value and a 12 percent coupon rate. Suppose a broker’s commission of Br.50 is imposed by brokers to buy or sell the bond. Assume further, that the discount rate (minimum


required rate of return) is 10 percent and the bond pays interest annually. What is


the price of the bond?


2. Project X requires an immediate investment of $150,000 and will generate net


cash inflows of $60,000 for the next three years. The project’s discount rate is 7%.


If net present value is used to appraise the project, should Project X be


undertaken?


1
Expert's answer
2022-03-07T10:43:33-0500

1.) Bond price.

Commission will increase face value by 50.

"Formula=C\times\frac{1-(1+r)^{-n}}{r}+\frac{FV}{(1+r)^n}"


Where; C is coupon payment="Coupon rate\times FV=.12\times1050=126"


FV=1,000+50=1,050


"Formula=126\times\frac{1-(1+0.1)^{-5}}{0.1}+\frac{1050}{(1+0.1)^5}=1,129.61"


Bond price=Br. 1,129.61


2.) Net Present Value.


"Year 0=-150,000\times\frac{1}{(1.07)^{0}}=-150,000"


"Year 1=60,000\times\frac{1}{(1.07)^{1}}=56,074.77"


"Year 2=60,000\times\frac{1}{(1.07)^{2}}=52,406.32"


"Year 3=60,000\times\frac{1}{(1.07)^{3}}=48,977.87"


To find NPV, sum all the above.

NPV=Year 0+ Year 1+ Year 2+ Year 2

NPV=-150,000+56,074.77+52,406.32+48,977.87=$7,458.96